Cosx+2cos(x+240°)

Simplify the equation


1
Expert's answer
2021-05-11T09:19:36-0400

cosx + 2cos(x+240"\degree") = cosx + 2(cosx * cos240"\degree" - sinx * sin240"\degree")

= cosx + 2(cosx * (-1/2) - sinx * (- "\sqrt{\smash[b]{3}}"/2))

= cosx + 2(-cosx/2 + "\sqrt{\smash[b]{3}}"sinx/2)

= cosx - cosx + "\sqrt{\smash[b]{3}}"sinx

= "\sqrt{\smash[b]{3}}"sinx


"\therefore" cosx + 2cos(x+240"\degree") = "\sqrt{\smash[b]{3}}"sinx